I have no idea how to do the following please help!
Suppose that a market research company finds that a price of p=$20, they would sell x=42 tiles each month. If they lower the price to p=$10, then more people would purchase teh tile, and they can expect to sell x=52 tiles in a month's time. Find the equation of the line for the demand equation. Write your answer in the form p=mx+b.
(write the equation using two points in the form(x,p)

Answer by stanbon(75887)

Suppose that a market research company finds that a price of p=$20, they would sell x=42 tiles each month. If they lower the price to p=$10, then more people would purchase teh tile, and they can expect to sell x=52 tiles in a month's time. Find the equation of the line for the demand equation. Write your answer in the form p=mx+b.
(write the equation using two points in the form (x,p)
-----------------------------------------------------------
You have two points: (42,20), (52,10)
---------------
slope = -10/10 = -1
-------
Form: p = mx+b
You know p, m, and x; solve for "b".
10 = -1*52+b
b = 62
EQUATION:
p(x) = -x + 62
--------------------
Cheers,
Stan H.